OpenStudy (anonymous):

Is 4 to 7 the same as 7 to 4? Explain why or why not

OpenStudy (anonymous):

will fan and medal

OpenStudy (mathstudent55):

Can you write the ratios as fractions?

OpenStudy (anonymous):

no I just says are they the same explain why or why not

OpenStudy (loser66):

You and your friend put the amount of money to buy a golden ring. You put 4 dollars, your friend put 7 dollars. However, your friend says, "4 to 7 is the same 7 to 4, so that you should put 7 dollars and I put 4 dollars. And we are best friend." What do you think?

OpenStudy (anonymous):

I don't know

OpenStudy (mathstudent55):

Ratios can be written in different ways such as: x:y x to y \(\dfrac{x}{y}\) All of the above expressions mean the ratio of x to y. When you write a ratio as a fraction, the rules of fractions apply. Let's write both of your ratios as fractions: \(\dfrac{4}{7} \) and \(\dfrac{7}{4} \) Since the rules of fractions apply to rations, the ratios 4 to 7 and 7 to 4 can only be equal if the fractions \(\dfrac{4}{7} \) and \(\dfrac{7}{4} \) are equal. Are these fractions equal?

OpenStudy (anonymous):

it aint fractions its part to part relationships

OpenStudy (mathstudent55):

A cake recipe states that you need to use 4 cups of flour to 7 eggs. The ratio of cups of flour to eggs is 4 to 7. You decide to use 7 cups of flour to 4 eggs. You used a ratio of cups of flour to eggs of 7 to 4. Your cake comes out tasting terrible and way too dry. Now you understand that 4 to 7 is different from 7 to 4.

OpenStudy (anonymous):

it aint a cake recipe

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

i have no idea what it would be

OpenStudy (anonymous):

the choices to choose from r yes or no is that better

OpenStudy (anonymous):

yes it does thank you I think that its a Yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome